The ability to taste the bitter compound phenylthiocarbamide (PTC) is an autosomal dominant trait. The inability to taste PTC is a recessive condition. In a sample of 500 people, 303 have the ability to taste PTC and 197 do not. Calculate the frequency of: the recessive allele, the dominant allele, the homozygous dominant genotype, the homozygous recessive genotype, the heterozygous genotype,
Answer #1

Answer:

Frequency of non-PTC = q^2 = 197/500 = 0.394

Frequency of non-PTC allele = q = SQRT of 0.394 = 0.628

Frequency of PTC allele = p = 1-0.628 = 0.372 (as p+q=1)

Therefore,

Frequency of recessive allele = 0.63

Frequency of the dominant allele = 0.37

Frequency of homozygous dominant genotype = p^2 = 0.372*0.372 = 0.139

Frequency of heterozygous genotype = pq = 2 * 0.372*0.628 = 0.467

Frequency of homozygous recessive genotype = q^2 = 0.394